# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 4.3 Two-step operator mixing
$\gamma_{\text{JMS}}^{ur}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&10eg_{s}^{2}m_{u}\over 9\pi^{2}&40eg_{s}^{2}m_{u}\over 27\pi^{2}\\ 0&-19g_{s}^{2}\over 3&g_{s}^{3}m_{u}\over 6\pi^{2}&-11g_{s}^{3}m_{u}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (111)
@@
$\gamma_{\text{JMS}}^{dr}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&-5eg_{s}^{2}m_{d}\over 9\pi^{2}&-20eg_{s}^{2}m_{d}\over 27\pi^{2}\\ 0&-19g_{s}^{2}\over 3&g_{s}^{3}m_{d}\over 6\pi^{2}&-11g_{s}^{3}m_{d}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (112)
@@ 4.4.4 ${C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}$ onto dipole
The leptonic ${C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}$ operator can break CP is $p\neq r$. In this case, we have two vectors, namely $\vec{\mathcal{C}}_{er}=\{{C_{\!\underset{pp}{e\gamma}}},{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}\}$ and $\vec{\mathcal{C}}_{ep}=\{{C_{\!\underset{rr}{e\gamma}}},{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}\}$ One example of the 2-loop Feynman diagrams governing this mixing is given in diagram (e) of Fig. 3. In the 2-loop ADM (104) one has to make replacements $q_{u}\to q_{e}$ and $q_{d}\to q_{e}$. However, since these are purely leptonic operators which exhibit only QED mixing we need to replace gluons in these diagrams with photons. So we need to take out the color factors in the original result. We found that the 2-loop $P$-type diagrams (which are the only ones allowed for the pure leptonic operators) shown in Fig. 4 of Ciuchini:1993ks has a universal color factor $C_{F}=(T^{A}T^{A})_{ij}=\delta_{ij}(N^{2}-1)/2N$ Therefore, we need to divide the elements of ADM by this overall color factor $C_{F}$. With all these adjustments, we arrived at the RG equation for ${C_{\!\underset{pp}{e\gamma}}}$
${16\pi^{2}\mu}({d{C_{\!\underset{pp}{e\gamma}}}/d\mu})={15e^{3}m_{e}Q_{e}\over 12\pi^{2}}{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}.$ (113)
@@
$[C^{S1,RR}_{uu}]_{2222}$ $-5\cdot 10^{-4}$ $-1\cdot 10^{-1}$ $-7\cdot 10^{-3}$
$[C^{S8,RR}_{uu}]_{2222}$ $1\cdot 10^{-4}$ $2\cdot 10^{-2}$ $-9\cdot 10^{-3}$
$[C^{S1,RR}_{uu}]_{1122}$ $1\cdot 10^{-2}$ $-8\cdot 10^{-5}$ $1\cdot 10^{-7}$ $2\cdot 10^{-5}$ $-2\cdot 10^{-7}$
$[C^{S8,RR}_{uu}]_{1122}$ $4\cdot 10^{-3}$ $10\cdot 10^{-4}$ $4\cdot 10^{-8}$ $8\cdot 10^{-6}$ $2\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2211}$ $1\cdot 10^{-2}$ $8\cdot 10^{-5}$ $3\cdot 10^{-7}$ $5\cdot 10^{-5}$ $-9\cdot 10^{-7}$
$[C^{S8,RR}_{ud}]_{2211}$ $4\cdot 10^{-3}$ $1\cdot 10^{-3}$ $8\cdot 10^{-8}$ $2\cdot 10^{-5}$ $-2\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2222}$ $1\cdot 10^{-2}$ $8\cdot 10^{-5}$ $6\cdot 10^{-6}$ $1\cdot 10^{-3}$ $-2\cdot 10^{-5}$
$[C^{S8,RR}_{ud}]_{2222}$ $4\cdot 10^{-3}$ $1\cdot 10^{-3}$ $2\cdot 10^{-6}$ $3\cdot 10^{-4}$ $-5\cdot 10^{-5}$
$[C^{S1,RR}_{uddu}]_{2112}$ $-6\cdot 10^{-2}$ $-6\cdot 10^{-3}$ $-1\cdot 10^{-6}$ $-3\cdot 10^{-4}$ $2\cdot 10^{-5}$
$[C^{S8,RR}_{uddu}]_{2112}$ $7\cdot 10^{-3}$ $-5\cdot 10^{-3}$ $1\cdot 10^{-7}$ $3\cdot 10^{-5}$ $9\cdot 10^{-6}$
$[C^{S1,RR}_{uddu}]_{2222}$ $-6\cdot 10^{-2}$ $-6\cdot 10^{-3}$ $-3\cdot 10^{-5}$ $-5\cdot 10^{-3}$ $4\cdot 10^{-4}$
$[C^{S8,RR}_{uddu}]_{2222}$ $7\cdot 10^{-3}$ $-5\cdot 10^{-3}$ $3\cdot 10^{-6}$ $6\cdot 10^{-4}$ $2\cdot 10^{-4}$
$[C^{S1,RR}_{uu}]_{1221}$ $-6\cdot 10^{-2}$ $-3\cdot 10^{-3}$ $-6\cdot 10^{-7}$ $-1\cdot 10^{-4}$ $-6\cdot 10^{-6}$
$[C^{S8,RR}_{uu}]_{1221}$ $7\cdot 10^{-3}$ $-5\cdot 10^{-3}$ $7\cdot 10^{-8}$ $1\cdot 10^{-5}$ $-10\cdot 10^{-6}$
$[C^{V1,LR}_{uddu}]_{2112}$ $-1\cdot 10^{-3}$ $-6\cdot 10^{-4}$ $-3\cdot 10^{-8}$ $-5\cdot 10^{-6}$ $-5\cdot 10^{-7}$
$[C^{V8,LR}_{uddu}]_{2112}$ $1\cdot 10^{-3}$ $-2\cdot 10^{-3}$ $3\cdot 10^{-8}$ $5\cdot 10^{-6}$ $5\cdot 10^{-6}$
$[C^{V1,LR}_{uddu}]_{2222}$ $-1\cdot 10^{-3}$ $-6\cdot 10^{-4}$ $-5\cdot 10^{-7}$ $-1\cdot 10^{-4}$ $-1\cdot 10^{-5}$
$[C^{V8,LR}_{uddu}]_{2222}$ $1\cdot 10^{-3}$ $-2\cdot 10^{-3}$ $6\cdot 10^{-7}$ $1\cdot 10^{-4}$ $1\cdot 10^{-4}$
$[C^{V1,LR}_{uu}]_{1221}$ $-1\cdot 10^{-3}$ $-2\cdot 10^{-3}$ $-1\cdot 10^{-8}$ $-3\cdot 10^{-6}$ $-3\cdot 10^{-6}$
$[C^{V8,LR}_{uu}]_{1221}$ $1\cdot 10^{-3}$ $-2\cdot 10^{-3}$ $1\cdot 10^{-8}$ $3\cdot 10^{-6}$ $-4\cdot 10^{-6}$
$[C_{u\gamma}]_{22}$ $-3\cdot 10^{-4}$ $-7\cdot 10^{-2}$ $9\cdot 10^{-1}$
$[C_{uG}]_{22}$ $7\cdot 10^{-3}$ $1\cdot 10^{0}$ $-9\cdot 10^{-2}$
$[C^{S,RR}_{eu}]_{1122}$ $-2\cdot 10^{-4}$ $6\cdot 10^{-8}$
$[C^{T,RR}_{eu}]_{1122}$ $7\cdot 10^{-2}$ $1\cdot 10^{-6}$ $-3\cdot 10^{-5}$
@@
Table 4: At the 2-loop level, the operator mixing of the charm quark operators in $4+1$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $4+1$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$ Entries below $10^{-8}$ are dropped.
All four-quark and dipole operators with bottom-quark in 5+3 WET
$\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$

### v2

@@ 4.3 Two-step operator mixing
$\gamma_{\text{JMS}}^{ur}=\begin{pmatrix}8g_{s}^{2}\over 3&{64}eg_{s}\over 9&10eg_{s}^{2}m_{u}\over 9\pi^{2}&40eg_{s}^{2}m_{u}\over 27\pi^{2}\\ 16eg_{s}\over 3&-19g_{s}^{2}\over 3&g_{s}^{3}m_{u}\over 6\pi^{2}&-11g_{s}^{3}m_{u}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (111)
@@
$\gamma_{\text{JMS}}^{dr}=\begin{pmatrix}8g_{s}^{2}\over 3&-32eg_{s}\over 9&-5eg_{s}^{2}m_{d}\over 9\pi^{2}&-20eg_{s}^{2}m_{d}\over 27\pi^{2}\\ -8eg_{s}\over 3&-19g_{s}^{2}\over 3&g_{s}^{3}m_{d}\over 6\pi^{2}&-11g_{s}^{3}m_{d}\over 72\pi^{2}\\ 0&0&0&-8g_{s}^{2}\over 3\\ 0&0&-12g_{s}^{2}&-14g_{s}^{2}\end{pmatrix}.$ (112)
@@
All four-quark and dipole operators with bottom-quark in 5+3 WET
$\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$ $\eta_{i}^{j}(\mu_{b},\mu_{\rm ew})$
$[C^{S1,RR}_{dd}]_{3333}$ $4.6\cdot 10^{-8}$ $9.8\cdot 10^{-8}$ $2.0\cdot 10^{-6}$ $-1.4\cdot 10^{-4}$ $-2.6\cdot 10^{-1}$ $1.2\cdot 10^{-2}$
$[C^{S8,RR}_{dd}]_{3333}$ $-2.0\cdot 10^{-8}$ $-4.1\cdot 10^{-7}$ $3.0\cdot 10^{-5}$ $5.5\cdot 10^{-2}$ $1.4\cdot 10^{-2}$
$[C^{S1,RR}_{dd}]_{1133}$ $2.6\cdot 10^{-2}$ $3.6\cdot 10^{-4}$ $1.3\cdot 10^{-8}$ $2.4\cdot 10^{-5}$ $1.4\cdot 10^{-7}$
$[C^{S8,RR}_{dd}]_{1133}$ $8.9\cdot 10^{-3}$ $-1.1\cdot 10^{-3}$ $8.1\cdot 10^{-6}$ $-1.2\cdot 10^{-6}$
$[C^{S1,RR}_{dd}]_{2233}$ $2.6\cdot 10^{-2}$ $3.6\cdot 10^{-4}$ $2.6\cdot 10^{-7}$ $4.8\cdot 10^{-4}$ $2.7\cdot 10^{-6}$
$[C^{S8,RR}_{dd}]_{2233}$ $8.9\cdot 10^{-3}$ $-1.1\cdot 10^{-3}$ $8.9\cdot 10^{-8}$ $1.6\cdot 10^{-4}$ $-2.4\cdot 10^{-5}$
$[C^{S1,RR}_{dd}]_{1331}$ $-1.7\cdot 10^{-1}$ $4.2\cdot 10^{-3}$ $-8.5\cdot 10^{-8}$ $-1.6\cdot 10^{-4}$ $5.9\cdot 10^{-6}$
$[C^{S8,RR}_{dd}]_{1331}$ $2.2\cdot 10^{-2}$ $8.2\cdot 10^{-3}$ $1.1\cdot 10^{-8}$ $2.0\cdot 10^{-5}$ $8.4\cdot 10^{-6}$
$[C^{S1,RR}_{dd}]_{2332}$ $-1.7\cdot 10^{-1}$ $4.2\cdot 10^{-3}$ $-1.7\cdot 10^{-6}$ $-3.2\cdot 10^{-3}$ $1.2\cdot 10^{-4}$
$[C^{S8,RR}_{dd}]_{2332}$ $2.2\cdot 10^{-2}$ $8.2\cdot 10^{-3}$ $2.2\cdot 10^{-7}$ $4.0\cdot 10^{-4}$ $1.7\cdot 10^{-4}$
$[C^{S1,RR}_{ud}]_{1133}$ $2.6\cdot 10^{-2}$ $-7.6\cdot 10^{-4}$ $1.1\cdot 10^{-5}$ $2.5\cdot 10^{-8}$
$[C^{S8,RR}_{dd}]_{1133}$ $8.9\cdot 10^{-3}$ $-1.1\cdot 10^{-3}$ $8.1\cdot 10^{-6}$ $-1.2\cdot 10^{-6}$
$[C^{S1,RR}_{uddu}]_{1331}$ $-1.7\cdot 10^{-1}$ $1.4\cdot 10^{-2}$ $-4.0\cdot 10^{-8}$ $-7.3\cdot 10^{-5}$ $-8.1\cdot 10^{-6}$
$[C^{S8,RR}_{uddu}]_{1331}$ $2.1\cdot 10^{-2}$ $6.9\cdot 10^{-3}$ $9.0\cdot 10^{-6}$ $-7.5\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2233}$ $-4.5\cdot 10^{-8}$ $9.9\cdot 10^{-5}$ $6.0\cdot 10^{-3}$ $1.3\cdot 10^{-5}$ $2.4\cdot 10^{-2}$ $-3.2\cdot 10^{-4}$
$[C^{S8,RR}_{ud}]_{2233}$ $-1.5\cdot 10^{-8}$ $3.2\cdot 10^{-5}$ $1.9\cdot 10^{-3}$ $6.5\cdot 10^{-4}$ $7.7\cdot 10^{-3}$ $-1.4\cdot 10^{-3}$
$[C^{S1,RR}_{uddu}]_{2332}$ $1.4\cdot 10^{-8}$ $2.9\cdot 10^{-7}$ $-6.5\cdot 10^{-4}$ $-3.9\cdot 10^{-2}$ $-4.3\cdot 10^{-3}$ $-1.6\cdot 10^{-1}$ $1.1\cdot 10^{-2}$
$[C^{S8,RR}_{uddu}]_{2332}$ $-3.6\cdot 10^{-8}$ $7.9\cdot 10^{-5}$ $4.8\cdot 10^{-3}$ $-4.0\cdot 10^{-3}$ $1.9\cdot 10^{-2}$ $7.5\cdot 10^{-3}$
$[C^{V1,LR}_{uddu}]_{1331}$ $-5.0\cdot 10^{-2}$ $1.1\cdot 10^{-3}$ $-1.7\cdot 10^{-6}$ $-2.3\cdot 10^{-6}$
$[C^{V8,LR}_{uddu}]_{1331}$ $1.3\cdot 10^{-2}$ $1.2\cdot 10^{-2}$ $2.6\cdot 10^{-6}$ $-5.5\cdot 10^{-6}$
$[C^{V1,LR}_{dd}]_{1331}$ $-4.9\cdot 10^{-2}$ $3.2\cdot 10^{-3}$ $-3.6\cdot 10^{-6}$ $4.1\cdot 10^{-6}$
$[C^{V8,LR}_{dd}]_{1331}$ $1.4\cdot 10^{-2}$ $1.2\cdot 10^{-2}$ $5.7\cdot 10^{-6}$ $5.6\cdot 10^{-6}$
$[C^{V1,LR}_{dd}]_{2332}$ $-4.9\cdot 10^{-2}$ $3.2\cdot 10^{-3}$ $-3.9\cdot 10^{-8}$ $-7.1\cdot 10^{-5}$ $8.1\cdot 10^{-5}$
$[C^{V8,LR}_{dd}]_{2332}$ $1.4\cdot 10^{-2}$ $1.2\cdot 10^{-2}$ $6.1\cdot 10^{-8}$ $1.1\cdot 10^{-4}$ $1.1\cdot 10^{-4}$
$[C^{V1,LR}_{uddu}]_{2332}$ $-1.8\cdot 10^{-4}$ $-8.6\cdot 10^{-4}$ $-1.2\cdot 10^{-3}$ $-4.5\cdot 10^{-2}$ $2.9\cdot 10^{-4}$
$[C^{V8,LR}_{uddu}]_{2332}$ $5.0\cdot 10^{-5}$ $1.3\cdot 10^{-3}$ $-2.7\cdot 10^{-3}$ $1.3\cdot 10^{-2}$ $1.2\cdot 10^{-2}$
$[C_{dG}]_{33}$ $-2.2\cdot 10^{-7}$ $-4.7\cdot 10^{-7}$ $-9.5\cdot 10^{-6}$ $-4.2\cdot 10^{-8}$ $6.9\cdot 10^{-4}$ $1.3\cdot 10^{0}$ $3.2\cdot 10^{-2}$
$[C_{d\gamma}]_{33}$ $-1.9\cdot 10^{-7}$ $1.3\cdot 10^{-5}$ $2.5\cdot 10^{-2}$ $9.0\cdot 10^{-1}$
Table 4: Operator mixing of the bottom quark operators in $5+3$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $5+3$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$, $\mu_{\rm ew}=91.1876$ GeV, $\mu_{\rm low}=2.0$ GeV . Entries below $10^{-8}$ are dropped. Note that for the charm and bottom dipoles, we stop running at $\mu=m_{b}=4.18$ GeV.
@@ 4.4.4 ${C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}$ onto dipole
The leptonic ${C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}$ operator can break CP is $p\neq r$. In this case, we have two vectors, namely $\vec{\mathcal{C}}_{er}=\{{C_{\!\underset{pp}{e\gamma}}},{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}\}$ and $\vec{\mathcal{C}}_{ep}=\{{C_{\!\underset{rr}{e\gamma}}},{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}\}$ One example of the 2-loop Feynman diagrams governing this mixing is given in diagram (e) of Fig. 3. In the 2-loop ADM (104) one has to make replacements $q_{u}\to q_{e}$ and $q_{d}\to q_{e}$. However, since these are purely leptonic operators which exhibit only QED mixing we need to replace gluons in these diagrams with photons. So we need to take out the color factors in the original result. We found that the 2-loop $P$-type diagrams (which are the only ones allowed for the pure leptonic operators) shown in Fig. 4 of Ciuchini:1993ks has a universal color factor $C_{F}=(T^{A}T^{A})_{ij}=\delta_{ij}(N^{2}-1)/2N$
Therefore, we need to divide the elements of ADM by this overall color factor $C_{F}$. With all these adjustments, we arrived at the RG equation for ${C_{\!\underset{pp}{e\gamma}}}$
${16\pi^{2}\mu}({d{C_{\!\underset{pp}{e\gamma}}}/d\mu})={15e^{3}m_{r}Q_{e}\over 12\pi^{2}}{C_{\!\underset{prrp}{ee}}^{{\tiny V,LR}}}.$ (113)
@@
$[C^{S1,RR}_{uu}]_{2222}$ $-3.7\cdot 10^{-4}$ $-1.0\cdot 10^{-1}$ $-7.0\cdot 10^{-3}$
$[C^{S8,RR}_{uu}]_{2222}$ $8.5\cdot 10^{-5}$ $2.3\cdot 10^{-2}$ $-8.5\cdot 10^{-3}$
$[C^{S1,RR}_{uu}]_{1122}$ $1.3\cdot 10^{-2}$ $-8.8\cdot 10^{-5}$ $9.2\cdot 10^{-8}$ $2.5\cdot 10^{-5}$ $-1.6\cdot 10^{-7}$
$[C^{S8,RR}_{uu}]_{1122}$ $4.4\cdot 10^{-3}$ $9.7\cdot 10^{-4}$ $3.0\cdot 10^{-8}$ $8.2\cdot 10^{-6}$ $1.8\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2211}$ $1.3\cdot 10^{-2}$ $8.4\cdot 10^{-5}$ $2.0\cdot 10^{-7}$ $5.4\cdot 10^{-5}$ $-9.5\cdot 10^{-7}$
$[C^{S8,RR}_{ud}]_{2211}$ $4.0\cdot 10^{-3}$ $1.1\cdot 10^{-3}$ $5.9\cdot 10^{-8}$ $1.6\cdot 10^{-5}$ $-2.4\cdot 10^{-6}$
$[C^{S1,RR}_{ud}]_{2222}$ $1.3\cdot 10^{-2}$ $8.4\cdot 10^{-5}$ $4.0\cdot 10^{-6}$ $1.1\cdot 10^{-3}$ $-1.9\cdot 10^{-5}$
$[C^{S8,RR}_{ud}]_{2222}$ $4.0\cdot 10^{-3}$ $1.1\cdot 10^{-3}$ $1.2\cdot 10^{-6}$ $3.3\cdot 10^{-4}$ $-4.9\cdot 10^{-5}$
$[C^{S1,RR}_{uddu}]_{2112}$ $-6.4\cdot 10^{-2}$ $-6.1\cdot 10^{-3}$ $-9.4\cdot 10^{-7}$ $-2.6\cdot 10^{-4}$ $1.7\cdot 10^{-5}$
$[C^{S8,RR}_{uddu}]_{2112}$ $6.9\cdot 10^{-3}$ $-4.9\cdot 10^{-3}$ $1.0\cdot 10^{-7}$ $2.7\cdot 10^{-5}$ $9.0\cdot 10^{-6}$
$[C^{S1,RR}_{uddu}]_{2222}$ $-6.4\cdot 10^{-2}$ $-6.1\cdot 10^{-3}$ $-1.9\cdot 10^{-5}$ $-5.2\cdot 10^{-3}$ $3.5\cdot 10^{-4}$
$[C^{S8,RR}_{uddu}]_{2222}$ $6.9\cdot 10^{-3}$ $-4.9\cdot 10^{-3}$ $2.0\cdot 10^{-6}$ $5.5\cdot 10^{-4}$ $1.8\cdot 10^{-4}$
$[C^{S1,RR}_{uu}]_{1221}$ $-6.4\cdot 10^{-2}$ $-3.4\cdot 10^{-3}$ $-4.4\cdot 10^{-7}$ $-1.2\cdot 10^{-4}$ $-6.3\cdot 10^{-6}$
$[C^{S8,RR}_{uu}]_{1221}$ $7.2\cdot 10^{-3}$ $-5.2\cdot 10^{-3}$ $4.9\cdot 10^{-8}$ $1.4\cdot 10^{-5}$ $-9.8\cdot 10^{-6}$
$[C^{V1,LR}_{uddu}]_{2112}$ $-1.3\cdot 10^{-2}$ $-2.3\cdot 10^{-3}$ $-1.7\cdot 10^{-8}$ $-4.8\cdot 10^{-6}$ $2.0\cdot 10^{-6}$
$[C^{V8,LR}_{uddu}]_{2112}$ $4.3\cdot 10^{-3}$ $-8.1\cdot 10^{-3}$ $3.4\cdot 10^{-8}$ $9.2\cdot 10^{-6}$ $9.2\cdot 10^{-6}$
$[C^{V1,LR}_{uddu}]_{2222}$ $-1.3\cdot 10^{-2}$ $-2.3\cdot 10^{-3}$ $-3.4\cdot 10^{-7}$ $-9.4\cdot 10^{-5}$ $4.0\cdot 10^{-5}$
$[C^{V8,LR}_{uddu}]_{2222}$ $4.3\cdot 10^{-3}$ $-8.1\cdot 10^{-3}$ $6.7\cdot 10^{-7}$ $1.8\cdot 10^{-4}$ $1.8\cdot 10^{-4}$
$[C^{V1,LR}_{uu}]_{1221}$ $-1.3\cdot 10^{-2}$ $-3.3\cdot 10^{-3}$ $-1.9\cdot 10^{-6}$ $-6.1\cdot 10^{-6}$
$[C^{V8,LR}_{uu}]_{1221}$ $4.5\cdot 10^{-3}$ $-8.0\cdot 10^{-3}$ $1.8\cdot 10^{-8}$ $5.0\cdot 10^{-6}$ $-8.3\cdot 10^{-6}$
$[C_{u\gamma}]_{22}$ $-4.9\cdot 10^{-4}$ $-1.4\cdot 10^{-1}$ $8.7\cdot 10^{-1}$
$[C_{uG}]_{22}$ $5.1\cdot 10^{-3}$ $1.4\cdot 10^{0}$ $-8.7\cdot 10^{-2}$
$[C^{S,RR}_{eu}]_{1122}$ $-1.5\cdot 10^{-4}$ $5.8\cdot 10^{-8}$
$[C^{T,RR}_{eu}]_{1122}$ $7.4\cdot 10^{-2}$ $1.5\cdot 10^{-6}$ $-3.0\cdot 10^{-5}$
@@
Table 5: Operator mixing of the charm quark operators in $4+1$ WET (1st column) onto dipole operators (2nd row) in $3+1$ and $4+1$ flavor WET is shown. Here the entries represent the quantity $\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$, $\mu_{\rm ew}=91.1876$ GeV, $\mu_{\rm low}=2.0$ GeV. Entries below $10^{-8}$ are dropped.
4.5 Numerical solution for the RGEs
In this section, we provide the numerical solution to the RGEs, which allows us to quantify the impact of such effects on the low-energy operators that contribute to EDMs at the tree level. As we have learned in the previous subsections, 1-loop and 2-loop RG running effects and threshold effects cause operator mixing leading to the expansion of operator basis for the EDMs.
All semileptonic operators in 5+3 WET
$\eta_{i}^{j}(\mu_{\rm low},\mu_{\rm ew})$ $\eta_{i}^{j}(\mu_{b},\mu_{\rm ew})$